# SplatCo: Structure-View Collaborative Gaussian Splatting for Detail-Preserving Rendering of Large-Scale Unbounded Scenes

### v3

@@ III-B Preliminary for 3D Gaussian Splatting
3D Gaussian Splatting [7] proposes a refined framework for scene representation, utilizing anisotropic 3D Gaussians as its core primitives. This method effectively integrates the differentiability inherent in volumetric representations with the computational efficiency of tile-based rasterization, achieving rapid rendering while preserving the capacity for gradient-based optimization.
@@
where $x^{\prime}\in\mathbb{R}^{2}$ represents a pixel coordinate, and $N$ denotes the set of depth-ordered 2D Gaussians contributing to that pixel. The differentiable nature of this rasterization pipeline facilitates end-to-end optimization of all Gaussian attributes—position, covariance, opacity, and color—using gradient-based methods, ensuring robustness and flexibility across diverse scene types.
3DGS excels at balancing rendering speed with visual quality, offering a compelling alternative to traditional volumetric rendering approaches. By harnessing hardware acceleration and avoiding per-pixel ray tracing, 3DGS provides an efficient pipeline well-suited for real-time applications while retaining the adaptability required for high-fidelity scene reconstruction.
III-C Cross-Structure Collaboration Module
@@
Therefore, we propose the Cross-Structure Collaboration Module (CSCM), which integrates the global perception of the tri-plane structure with the local sensitivity of the context grid structure. This approach facilitates more accurate feature extraction for each Gaussian point within the overall point cloud, significantly improving the quality of rendering details in subsequent novel view synthesis. Inspired by Feature Pyramid Networks (FPN) [64], we divide the overall training process into three levels. During training, features are progressively compensated from coarse to fine, aiming to reduce feature misalignment caused by the size of the tri-plane feature map and grid resolution. Specifically, the CSCM consists of three components: tri-plane feature acquisition, context interpolation, and cross-structure hierarchical compensation.
Tri-plane Feature Acquisition. Tri-Plane is a static 3D instance of the $k$-planes factorization [65], which represents the 3D space using three $m$-channel feature plane maps $\mathbf{F}\in\mathbb{R}^{w\times h\times m}$, each corresponding to one of the axis-aligned planes: $xy$, xzxz, and $yz$, with spatial resolution $w\times h$. To obtain fine-grained tri-plane features for a queried 3D point $\mathbf{p}$, we first apply a tri-plane attention module to generate attention-enhanced tri-plane features $\mathbf{F}^{TA}$. This module captures relative features across all three planes from the global tri-plane context, thereby enhancing the global perception capability compared to traditional tri-plane representations that focus solely on individual planes. The formulation is as follows:
@@
where $i$ denotes the axis-aligned planes: $xy$, xzxz, and $yz$, and $\text{TA}(\cdot)$ represents the tri-plane attention operation. The operator $||_{j\in\{xy,xz,yz\}}(\cdot)$ indicates concatenating feature planes along the channel dimension in the order of $xy$, xzxz, and $yz$. $\text{SA}(\cdot)$ and $\text{CA}(\cdot)$ denote the spatial attention and channel attention operations [66] applied to the feature maps, respectively. $\rho_{i}(\cdot)$ corresponds to the operation of extracting the tri-plane attention feature map for the dimension indexed by $i$ from the previously concatenated feature maps.
@@
where $\pi_{i}(\cdot)$ is the projection function, and $\psi(\cdot)$ denotes bi-linear interpolation on the feature map. $\phi_{t}(\cdot)$ represents the tri-plane feature decoder. We first concatenate the interpolated original features from both the original and attention-enhanced feature maps of the same orientation. Then, the features from different tri-plane orientations are sequentially concatenated along the channel dimension. Finally, the concatenated features are passed through the decoder to obtain the tri-plane feature $f_{t}$.
Context Interpolation. In addition to leveraging tri-plane features for global scene perception, we aim to uncover the local intrinsic relationships between the 3D Gaussians that construct the scene. To achieve this, we employ grid-based structures to obtain contextual interpolation features for each Gaussian point, which are typically compact, well-organized, and capable of establishing spatial connections among the Gaussians.
To obtain the contextual interpolation information for the current query Gaussian point $p$, we first create a grid structure for the current point cloud based on the given resolution $\alpha$. As shown in Fig. 2(b), the vertices of the cube with blue contours represent the context Gaussians for interpolation of the query Gaussian. Then, based on the distance between the query Gaussian point $p$ and the vertices of the corresponding grid, we use the features $f_{g}$ of the vertex Gaussians to perform a weighted aggregation. Finally, the decoder $\phi_{c}$ is used to decode and obtain the contextual feature $f_{c}$ for the current point:
@@
Cross-Structure Hierarchical Compensation. Once we have obtained the global features $f_{t}$ of the query point $p$ from the tri-plane structure and the local features $f_{c}$ from the contextual grid interpolation, we propose a cross-structure compensation mechanism to fully leverage the advantages of both types of features. This mechanism results in the refined feature $f_{h}$, which captures both global and local point cloud relationships:
@@
Additionally, considering that the refinement of scene representation by both structural features is limited to some extent by the resolution of the feature maps and the grid resolution, we introduce a hierarchical compensation mechanism based on the cross-structure refined features. We divide the overall training process into three levels, progressively increasing the resolution of the tri-plane feature maps and utilizing finer contextual grids as the training progresses. The refined features $f_{h}$ from each level are then merged:
@@
III-D Cross-View Assisted Training Strategy
Although we have previously derived cross-structure collaborative features that effectively capture the Gaussian point distribution, the inherent reliance on single-view image-based optimization in the original 3DGS method exacerbates the overfitting of Gaussians to the training viewpoint. As illustrated in the left side of Fig. 2 (c), the Gaussian ellipsoids compensate for the current training viewpoint by being distributed extremely close to the camera position. However, they do not lie in positions that are realistically generalizable to new viewpoints, leading to significantly better rendering quality for the training viewpoint compared to the test viewpoint, thus causing overfitting. Moreover, in large-scale scenes, the overfitting issue becomes more pronounced due to the limited overlap between images from different viewpoints.
To address this issue, we integrate the explicit multi-view regularization strategy proposed in MVGS [13] into our cross-view assisted training strategy. Specifically, we first employ a Multi-View Regulated Learning method from MVGS  [13] that leverages multiple views for collaborative optimization. Additionally, we also adopt the Multi-View Augmented Densification in MVGS [13] to improve point distribution by incorporating information from spatially distant views, which is particularly advantageous in large-scale scenes. Finally, a Cross-View Consistent Loss is introduced to enforce coherence between adjacent viewpoints and mitigate overfitting of Gaussians.
Multi-View Regulated Learning [13]. We extend the original single-view iterative training to a multi-view stereo approach that optimizes 3D Gaussians under multi-view supervision. In each iteration, we sample $M$ pairs of images and camera parameters, ensuring they are distinct from one another. We adopt the same multi-view regularization as in MVGS [13]:
$\frac{\partial\mathcal{L}}{\partial\{G\}}=\frac{\partial\mathcal{L}(G_{1},E_{1},K_{1})}{\partial G_{1}}+\cdots+\frac{\partial\mathcal{L}(G_{M},E_{M},K_{M})}{\partial G_{M}},$ (12)
where $\mathcal{L}$ represents the loss at the current iteration, $G=\{G_{1},G_{2},\dots,G_{M}\}$ denotes the set of 3D Gaussians, with each $G_{M}$ being affected by large gradients from each view during multi-view training, and $E_{M}$ and $K_{M}$ are the extrinsic and intrinsic parameters of the camera, respectively.
The key distinction from the original 3DGS loss is that our approach incorporates a multi-view gradient constraint to optimize the 3D Gaussians. This allows each Gaussian kernel $g_{i}$ to be adjusted based on multi-view information, effectively reducing overfitting to individual views. Additionally, the multi-view constraint helps the 3D Gaussians learn view-dependent features, such as reflections (illustrated in the left part of Fig. 2(c)), enabling improved performance in novel view synthesis for reflective scenes.
Multi-View Augmented Densification [13]. When optimizing across multiple viewpoints, particularly in large-scale scenes, selecting views that are too far apart is common. This results in limited overlap between perspectives, which increases the risk of overfitting Gaussian ellipsoids. To address this, we use the same distance-aware multi-view densification in MVGS [13]. This process involves normalizing the camera translations of sampled views to fit within a unit sphere, making our method adaptable to various scenes. Then it computes the relative translation distances $\{r_{i}|i=1,2,\dots,M\}$ between the cameras. The adaptive densification threshold with a self-adaptive threshold $\hat{\beta}$ in MVGS [13] is as follows:
$\hat{\beta}=\frac{\beta}{2}H\left(\frac{r_{i}}{\tau}-1\right)+\beta\left(1-H\left(\frac{r_{i}}{\tau}-1\right)\right),$ (13)
where $H(\cdot)$ is the Heaviside function, returning 1 if the input is greater than or equal to 0. $\beta$ is the threshold for the original 3DGS densification, and $\tau$ is a predefined hyperparameter.
This approach ensures that when discrepancies between views increase, the extent of 3D Gaussian densification is also enhanced.
Cross-View Consistent Loss. To enhance the continuity of scene structures and reduce overfitting of Gaussians, we first select adjacent views from the $M$ viewpoints, using the Structural Similarity Index Measure (SSIM) between images as a selection criterion, with the condition that the SSIM value is greater than 0.6. Next, we compute the Cross-view Consistent Loss $L_{\text{CVC}}$ between two images and obtain a mask $M_{f}$ to remove overfitted Gaussians:
$L_{\text{CVC}}=SSIM(I_{i}^{gt},I_{j}^{gt})*L_{1}(I_{i}^{gt}-I_{i},I_{j}^{gt}-I_{j}),$ (14)
$M_{f}(\mathbf{p})=\begin{cases}1,&\text{if}\,\alpha(\mathbf{p})=True\\ 0,&\text{otherwise}\end{cases}$ (15)
$\alpha(\mathbf{p})=d_{I}(\mathbf{p})<0.01\land(d_{C}(\mathbf{p})<0.5\lor d_{O}(\mathbf{p})>3\sigma).$ (16)
Here, $M_{f}(\mathbf{p})=1$ indicates that the Gaussian point $\mathbf{p}$ is considered an overfitted Gaussian and should be pruned. $d_{I}$ represents the distance from the Gaussian point to the camera’s viewpoint ray, and when $d_{I}<0.01$, we assume that the point is intersected by the viewpoint ray. $d_{C}$ is the distance from the Gaussian point to the camera position, and $d_{O}$ denotes the distance from the Gaussian point to the average position of the point cloud. If $d_{C}<0.5$ or $d_{O}>3\sigma$ where $\sigma$ is the point cloud’s variance, we regard the Gaussian as a visual overfitting point in the training view that neglects scene structure consistency, and it will be pruned. In this way, we optimize the distribution of Gaussians during training, effectively reducing the risk of overfitting.
TABLE I: Quantitative Comparison on Mill-19 and MatrixCity datasets. The ’-’ symbol represents Mega-NeRF [67] and Switch-NeRF [68] were not evaluated on MatrixCity. The best, second-best, and third-best results are marked in red, orange, and yellow, respectively.
MatrixCity Rubble Building
[…]
Mega-NeRF [CVPR22] [67]
0.553 24.06 0.516 0.547 20.93 0.504
Switch-NeRF [ICLR22] [68]
0.562 24.31 0.496 0.579 21.54 0.474
3DGS [ToG23] [7] 0.737 23.79 0.384 0.746 24.30 0.324 0.676 20.11 0.372
GaMeS [arXiv24] [69] 0.773 24.64 0.367 0.767 24.64 0.311 0.698 20.97 0.329
Compact3DGS [CVPR24] [70] 0.795 26.03 0.370 0.719 23.01 0.373 0.645 20.00 0.425
Scaffold-GS [CVPR24] [71] 0.842 27.25 0.290 0.736 23.71 0.346 0.711 21.55 0.328
Pixel-GS [ECCV24] [72] 0.814 25.60 0.324 0.771 24.67 0.308 0.700 21.00 0.352
MVGS [arxiv24] [13] 0.836 26.92 0.271 0.772 25.82 0.268 0.731 22.46 0.272
Ours 0.855 28.05 0.249 0.794 26.90 0.280 0.759 23.03 0.270
III-E Update Gaussian Attributes &\& Rasterization
To fully leverage the features previously acquired, we employ the method of neural Gaussians [71] obtained earlier.
@@ III-E Update Gaussian Attributes &\& Rasterization
We integrate Gaussian positions, camera poses, and hierarchical features through multiple MLPs to update Gaussian attributes such as position $\mu\in\mathbb{R}^{3}$, opacity $\alpha\in\mathbb{R}$, quaternion $q\in\mathbb{R}^{4}$, scaling $s\in\mathbb{R}^{3}$, and color $c\in\mathbb{R}^{3}$. For each visible anchor, $k$ Gaussians are generated and their respective attributes are computed. The position of the Gaussian primitives corresponding to an anchor at $x^{a}$ is derived as follows:
@@
After determining the parameters for each Gaussian, we use rasterization methods as outlined in 3DGS[7] to generate rendered images from the respective viewpoints. The network parameters are then optimized according to the loss functions detailed below.
IV Loss Function
In summary, our loss function consists of three components: 1) rendering pixel loss $L_{\text{pixel}}$, 2) Structural Similarity Index Measure (SSIM) loss $L_{\text{SSIM}}$ [73], 3) fine-grained Gaussian regularization loss $L_{\text{fine}}$, 4) cross-view consistent loss $L_{\text{CVC}}$.
IV-A Rendering Loss

### v4

@@ 3.2 Preliminary for 3D Gaussian Splatting
3DGS [relate_3dgs] proposes a new paradigm for scene representation, utilizing anisotropic 3D Gaussians as its core primitives. This method effectively integrates the differentiability inherent in volumetric representations with the computational efficiency of tile-based rasterization, achieving rapid rendering while preserving the capacity for gradient-based optimization.
@@
where $x^{\prime}\in\mathbb{R}^{2}$ represents a pixel coordinate, and $N$ denotes the set of depth-ordered 2D Gaussians contributing to that pixel. The differentiable nature of this rasterization pipeline facilitates end-to-end optimization of all Gaussian attributes—position, covariance, opacity, and color—using gradient-based methods, offering a compelling alternative to traditional volumetric rendering approaches.
3.3 Cross-Structure Collaboration Module
@@
Therefore, we propose the CSCM, which integrates the global perception of the tri-plane structure with the local sensitivity of the context grid structure. This approach facilitates more accurate feature extraction for each Gaussian point within the overall point cloud, significantly improving the quality of rendering details in subsequent novel view synthesis. Inspired by Feature Pyramid Networks (FPN) [method_fpn], we divide the overall training process into three levels. During training, features are progressively compensated from coarse to fine, aiming to reduce feature misalignment caused by the size of the tri-plane feature map and grid resolution. Specifically, the CSCM consists of three components: tri-plane feature acquisition, context interpolation, and cross-structure hierarchical compensation.
Tri-plane Feature Acquisition. Tri-Plane is a static 3D instance of the $k$-planes factorization [method_kplanes], which represents the 3D space using three $m$-channel feature plane maps $\mathbf{F}\in\mathbb{R}^{w\times h\times m}$, each corresponding to one of the axis-aligned planes: $xy$, xzxz, and $yz$, with spatial resolution $w\times h$. To obtain fine-grained tri-plane features for a queried 3D point $\mathbf{p}$, we first apply a tri-plane attention module to generate attention-enhanced tri-plane features $\mathbf{F}^{TA}$. This module captures relative features across all three planes from the global tri-plane context, thereby enhancing the global perception capability compared to traditional tri-plane representations that focus solely on individual planes. The formulation is as follows:
@@
where $i$ denotes the axis-aligned planes: $xy$, xzxz, and $yz$, and $\text{TA}(\cdot)$ represents the tri-plane attention operation. The operator $||_{j\in\{xy,xz,yz\}}(\cdot)$ indicates concatenating feature planes along the channel dimension in the order of $xy$, xzxz, and $yz$. $\text{SA}(\cdot)$ and $\text{CA}(\cdot)$ denote the spatial attention and channel attention operations [method_cbam] applied to the feature maps, respectively. $\rho_{i}(\cdot)$ corresponds to the operation of extracting the tri-plane attention feature map for the dimension indexed by $i$ from the previously concatenated feature maps.
@@
where $\pi_{i}(\cdot)$ is the projection function, and $\psi(\cdot)$ denotes bi-linear interpolation on the feature map. $\phi_{t}(\cdot)$ represents the tri-plane feature decoder.
Finally, We concatenate the interpolated original features from both the original and attention-enhanced feature maps of the same orientation. The features from different tri-plane orientations are sequentially concatenated along the channel dimension. The concatenated features are passed through the decoder to obtain the tri-plane feature $f_{t}$.
Context Interpolation. In addition to leveraging tri-plane features for global scene perception, we also exploit the local intrinsic relationships between the three-dimensional Gaussians to enrich the scene representation. To achieve this, we employ grid-based structures to obtain contextual interpolation features for each Gaussian point, which are typically compact, well-organized, and capable of establishing spatial connections among the Gaussians.
Specifically, for the current query Gaussian point $p$, we first create a grid structure for the current point cloud based on the given resolution $\alpha$. As shown in Fig. 2(b), the vertices of the cube with blue contours represent the context Gaussians for interpolation of the query Gaussian. Then, based on the distance between the query Gaussian point $p$ and the vertices of the corresponding grid, we use the features $f_{g}$ of the vertex Gaussians to perform a weighted aggregation. Finally, the decoder $\phi_{c}$ is used to decode and obtain the contextual feature $f_{c}$ for the current point:
@@
Cross-Structure Hierarchical Compensation. Once we have separately obtained the global features $f_{t}$ of the query point $p$ from the tri-plane structure and the local features $f_{c}$ from the contextual grid interpolation, we propose a cross-structure compensation mechanism to fully fuse the advantages of both types of features. This mechanism results in the refined Hierarchical Dense Embedded (HDE) feature $f_{h}$, which captures both global and local point cloud relationships:
@@
Additionally, considering that the refinement of scene representation by both structural features is limited to some extent by the resolution of the feature maps and the grid resolution, we introduce a hierarchical compensation mechanism based on the cross-structure refined features. We divide the overall training process into three levels, progressively increasing the resolution of the tri-plane feature maps and utilizing finer contextual grids as the training progresses. The refined HDE features $f_{h}$ from each level are then merged:
@@
3.4 Cross-View Pruning Mechanism
Although we have previously derived cross-structure collaborative HDE features $f_{h}$ that effectively capture the Gaussian point geometric structure information, the inherent reliance on single-view image-based optimization in the original 3DGS method exacerbates the overfitting of Gaussians to the training viewpoint.
Encouragingly, MVGS [mvgs] utilizes multi-view information for Gaussian densification, but this approach inevitably introduces numerous redundancies and escalates computational overhead. Since these Gaussians tend to overfit the current training viewpoint by clustering extremely close to the camera plane, they lack geometric validity for novel viewpoints despite minimizing training error. Moreover, this issue is further exacerbated in large-scale scenes where the limited overlap between viewpoints provides insufficient geometric constraints. To mitigate this issue, we introduce the CVPM, which incorporates the formulation of a Cross-View Consistency Loss and a subsequent Pruning Procedure to ensure the validity of stored Gaussians, thereby enhancing their geometric representativeness while simultaneously avoiding Gaussian rendering artifacts.
Cross-View Consistent Loss. To reinforce structural continuity across the scene and mitigate Gaussian overfitting, following  [mvgs], we first select adjacent view pairs from the set of $M$ views. The selection is governed by the Structural Similarity Index Measure (SSIM), where we retain pairs exhibiting an SSIM value greater than 0.6 to ensure sufficient geometric overlap. Subsequently, we compute the Cross-View Consistent Loss (CVCL), $L_{\text{CVC}}$, which penalizes discrepancies between the residuals of paired views, weighted by their structural similarity:
$L_{\text{CVC}}=\text{SSIM}(I_{i}^{gt},I_{j}^{gt})\cdot L_{1}\left((I_{i}^{gt}-I_{i}),(I_{j}^{gt}-I_{j})\right),$ (12)
where $I^{gt}$ and $I$ denote the ground truth and rendered images, respectively. This loss enforces that regions with high structural similarity in ground truth should exhibit consistent error patterns in reconstruction, thereby encouraging global coherence.
Pruning Procedure. Building upon the consistency constraints, we further implement a geometric pruning strategy to reduce Gaussian redundancies. We define a binary mask $M_{f}$ to identify and remove overfitted Gaussians:
$M_{f}(\mathbf{p})=\begin{cases}1,&\text{if}\,\,\alpha(\mathbf{p})\text{ is True},\\ 0,&\text{otherwise},\end{cases}$ (13)
where the pruning condition $\alpha(\mathbf{p})$ is determined by a combination of ray intersection and geometric plausibility criteria:
$\alpha(\mathbf{p})=\left(d_{I}(\mathbf{p})<0.01\right)\land\left(d_{C}(\mathbf{p})<0.5\lor d_{O}(\mathbf{p})>3\sigma\right).$ (14)
Here, $M_{f}(\mathbf{p})=1$ signifies that a Gaussian at position $\mathbf{p}$ is classified as an artifact and subsequently pruned. The term $d_{I}(\mathbf{p})$ represents the perpendicular distance from the Gaussian to the camera ray; a value $d_{I}<0.01$ implies the point lies essentially on the ray path. To filter out structurally invalid points, we evaluate two additional metrics: $d_{C}(\mathbf{p})$, the distance to the camera center, and $d_{O}(\mathbf{p})$, the distance to the point cloud centroid. A Gaussian is deemed a visual overfitting artifact that neglects structural consistency if it is either anomalously close to the camera ($d_{C}<0.5$) or constitutes a statistical outlier ($d_{O}>3\sigma$, where $\sigma$ is the spatial variance of the point cloud). Through this mechanism, we refine the Gaussian distribution during training, effectively curbing the risk of overfitting, while simultaneously reducing memory storage.
3.5 Structure-View Co-learning
Current 3DGS approaches universally employ view-based loss functions to drive the optimization of Gaussian parameters via gradient descent. The standard formulation for the view-based gradient is expressed as:
$\nabla\mathcal{L}=\sum_{i=1}^{M}\nabla_{\mathcal{V}_{i}}\mathcal{L}_{i},$ (15)
where $M$ denotes the number of training views, and $\nabla_{\mathcal{V}_{i}}\mathcal{L}_{i}$ represents the gradient derived from the $i$-th view.
However, in the context of large-scale scenes, this paradigm faces significant challenges. Characterized by extensive spatial coverage and sparse overlap between adjacent viewing sequences, view-dependent signals alone are insufficient to explicitly capture the holistic scene structure. Furthermore, approaches relying exclusively on view-based gradient backpropagation, whether from single[relate_3dgs] or multiple perspectives[mvgs], neglect this intrinsic structural information. Consequently, this renders the reconstruction highly susceptible to overfitting, typically manifesting as severe visual artifacts and blurred geometric structures.
To address these limitations overlooked by prior work, we naturally propose the SVC module. Moving beyond reliance solely on view-based gradients, our approach also leverages the proposed hierarchical context and multi-resolution tri-plane modules to extract structural gradient information for each viewpoint. The total gradient for joint optimization is formulated as:
$\nabla\mathcal{L}=\sum_{i=1}^{M}(\nabla_{\mathcal{V}_{i}}\mathcal{L}_{i}+\nabla_{\mathcal{S}}\mathcal{L}_{i}),$ (16)
where $\nabla_{\mathcal{S}}\mathcal{L}_{i}$ represents the structural gradient derived from our modules. By aggregating these intrinsic scene structural gradients with view gradients, the SVC module effectively optimizes the geometric position and appearance attributes of the Gaussians. It redirects the update of Gaussian parameters from a trajectory highly susceptible to overfitting towards one that is robustly constrained by structural guidance. This paradigm shift holds profound significance for advancing both the foundational 3DGS architecture and robust Novel View Synthesis in large-scale scenarios.
3.6 Update Gaussian Attributes &\& Rasterization
@@
To fully leverage the features previously acquired, we also employ the method of Neural Gaussians[relate_scaffoldgs] obtained earlier. Specifically, we integrate Gaussian positions, camera poses, and hierarchical features through multiple MLPs to update Gaussian attributes. For each visible anchor, $k$ Gaussians are generated and their respective attributes are computed. The position of the Gaussian primitives corresponding to an anchor at $x^{a}$ is derived as follows:
@@
After determining the parameters for each Gaussian, we use rasterization methods as outlined in 3DGS[relate_3dgs] to generate rendered images from the respective viewpoints. The network parameters are then optimized according to the loss functions.
4 Loss Function
In summary, our loss function consists of three components: 1) rendering pixel loss $L_{\text{pixel}}$, 2) Structural Similarity Index Measure (SSIM) loss $L_{\text{SSIM}}$ [exp_sim], 3) fine-grained Gaussian regularization loss $L_{\text{fine}}$, 4) cross-view consistent loss $L_{\text{CVC}}$.
4.1 Rendering Loss